OpenStudy (anonymous):

Find the indefinite integral of x/(x^4+2x^2+2) dx

OpenStudy (anonymous):

\[\int \dfrac{x}{x^4+2x^2+2} dx\] Let u = √x ... ↔ ... u² = x. 2u du = dx \[\int \dfrac{u^2}{u^2+2u+2} (2u\; du)\\ 2\int \dfrac{u^3}{u^2+2u+2} du\] Does that seem more doable?